A car of length 5 m is being driven parallel alongside a 100 m long train at the same

speed. The front of the car is 10 m behind the rear of the train. At the instant the train

decelerates at 0.4 ms2 the car accelerates at 0,15 ms2. How long will it take for the rear

of the car to pass the front of the train?


1
Expert's answer
2021-08-10T05:54:58-0400
"5+10+100+vt-0.5(0.4)t^2=vt+0.5(0.15)t^2\\t=20.45\ s"